OpenStudy (anonymous):

Let f(x) = xe^−x2/2 . i. Find all local maxima and minima of f(x). ii. Determine the intervals on which f(x) is decreasing. iii. Determine the intervals on which f(x) is increasing. Any assistance is appreciated, Kind regards

OpenStudy (anonymous):

Well, can you find the derivative first?

OpenStudy (anonymous):

1 moment

OpenStudy (anonymous):

the function is a little ambiguous... is it \(\large f(x)=xe^{-\frac{x^2}{2}} \) ????

OpenStudy (anonymous):

Agreed.

OpenStudy (anonymous):

yep...thats it

OpenStudy (anonymous):

where is the ambiguity?

OpenStudy (anonymous):

You didn't use any parentheses or notation. The function could have gone a number of ways.

OpenStudy (anonymous):

sorry...the one that dpa posted is correct...i'll be more careful next time

OpenStudy (anonymous):

so I factor out the x

OpenStudy (anonymous):

sorry...I should use the product rule

OpenStudy (anonymous):

Anyway to cut a long story short

OpenStudy (anonymous):

the derivative is y=(1-x^2).e((-x^2)/2)

OpenStudy (anonymous):

Do you know how to find the minima and maxima using the derivative?

OpenStudy (anonymous):

Could you advise me please . Do I need a 2nd derivative as well?

OpenStudy (anonymous):

You don't have to, no. But you know that there is a minima or maxima where the derivative is 0. Here, there are only two.

OpenStudy (anonymous):

\[y''=x(x^2-3)e-^{x ^{2}/2}\]

OpenStudy (anonymous):

okay..how do you tell that there are only two ?

OpenStudy (anonymous):

okay I plugged the function into my calulator and I get two 0s for the function. One where x is -1, the other where x is 1. So these two points are the maximum and minimium points...Is that correct....

OpenStudy (anonymous):

Sorry, yes that is correct. You'll see that the x^2 - 1 term factors into (x-1)(x+1). Here the exponential term has no zero, it will never cross the x-axis.

OpenStudy (anonymous):

That is, the term with 'e', in it.

OpenStudy (anonymous):

Thank you for that tip....! How do i determine where the intervals of f(x) is increasing or decreasing?

OpenStudy (anonymous):

You can set up some simple test values. Here, you can check the -1. Try plugging -2 and 0 into the original (non-differentiated) equation. For the 1, try 0 (which you already have), and 2.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):